OpenStudy (anonymous):

Sandi bought 6 yards of material. She wants to cut it into 2 pieces so that the difference between the lengths are 1.5 yards. What should be the length of each piece?

jimthompson5910 (jim_thompson5910):

Let x and y be the two lengths of the two pieces

jimthompson5910 (jim_thompson5910):

"Sandi bought 6 yards of material" means x+y = 6

jimthompson5910 (jim_thompson5910):

"She wants to cut it into 2 pieces so that the difference between the lengths are 1.5 yards" so x-y = 1.5 where x is implied to be the longer piece

jimthompson5910 (jim_thompson5910):

So you have the system x+y = 6 x-y = 1.5

OpenStudy (anonymous):

how do you solve that @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Add the equations

jimthompson5910 (jim_thompson5910):

x+y = 6 x-y = 1.5 ----------- 2x = 7.5 2x = 7.5 x = 7.5/2 x = 3.75 So the first piece is 3.75 yards

OpenStudy (anonymous):

got iit! THANKS SO MUCH

jimthompson5910 (jim_thompson5910):

x+y = 6 3.75 + y = 6 y = 6 - 3.75 y = 2.25 and the second piece is 2.25 yards